OpenStudy (wwe123):

the roots of the equation x²+px+8=0 are α and β.obtain the values of p,if

OpenStudy (wwe123):

\[(i) \alpha=\beta^{2}\]

OpenStudy (wwe123):

\[(ii) \alpha-\beta=2\]

OpenStudy (anonymous):

for first one\[\alpha \beta=\beta^3=8\]so u can evaluate roots easily do same thing for second one

OpenStudy (wwe123):

how \[ \alpha \beta=\beta^{2}\] @mukushla

OpenStudy (anonymous):

that is 3 beta^3

OpenStudy (wwe123):

3 beta^3 ??

OpenStudy (anonymous):

i mean the power is 3

OpenStudy (anonymous):

ab=b^3

OpenStudy (wwe123):

by how ab=b^3

OpenStudy (anonymous):

u have a=b^2 right ?

OpenStudy (wwe123):

yes

OpenStudy (anonymous):

so ab=(b^2)b=b^3

OpenStudy (wwe123):

but it become ab^2=b^3